HETAV PATEL

Grade 11,

evaluate the following integral ∫ 1 – sin x / x + cos x dx

evaluate the following integral
∫ 1 – sin x / x + cos x dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello student,
Please find the answer to your question below

I = \int \frac{1-sinx}{x+cosx}dx
x+cosx = t
(1-sinx)dx = dt
I = \int \frac{1}{t}dt
I = log(t) + c
I = log(x+cosx) + c